y = x^2 + (m+5)x+(5m+1) a perfect square trinomial.
Answer by ReadingBoosters(3246)

Starting with
y = x^2 + (m+5)x + (5m+1)
Distribute the x
y = x^2 + mx + 5x + 5m + 1
y = x^2 + 5x + mx +5m + 1
y = x(x+5)+m(x+5)+1
y = (x+m)(x+5) + 1
(x+m)(x+5)=-1
x+m=-1 so x= -1-m
x+5=-1 so x= -1-5 = -6